# Improving Coverage in Combined Prediction Sets with Weighted p-values

### v1

@@ Abstract
Conformal prediction quantifies the uncertainty of machine learning models by augmenting point predictions with valid prediction sets, assuming exchangeability. For complex scenarios involving multiple trials, models, or data sources, conformal prediction sets can be aggregated to create a prediction set that captures the overall uncertainty, often improving precision. However, aggregating multiple prediction sets with individual $1-\alpha$ coverage inevitably weakens the overall guarantee, typically resulting in $1-2\alpha$ worst-case coverage. In this work, we propose a framework for the weighted aggregation of prediction sets, where weights are assigned to each prediction set based on their contribution. Our framework offers flexible control over how the sets are aggregated, achieving tighter coverage bounds that interpolate between the $1-2\alpha$ guarantee of the combined models and the $1-\alpha$ guarantee of an individual model depending on the distribution of weights. We extend our framework to data-dependent weights, and we derive a general procedure for data-dependent weight aggregation that maintains finite-sample validity. We demonstrate the effectiveness of our methods through experiments on synthetic and real data in the mixture-of-experts setting, and we show that aggregation with data-dependent weights provides a form of adaptive coverage.
@@ 1 Introduction
In recent years, machine learning models have achieved remarkable accuracy across a variety of predictive tasks [1, 2, 3]. Understanding the uncertainty associated with each prediction is essential to decision-making in real-world scenarios; however, the black box nature of many machine learning models hinders their deployment in safety-critical applications such as medical diagnosis [4, 5, 6, 7], industrial control systems [8, 9], and extreme weather forecasting [10, 11, 12]. Conformal prediction [13] has emerged as a popular wrapper method around machine learning models because it provides a statistically valid quantification of uncertainty. In particular, it transforms point predictions to prediction sets with finite-sample coverage guarantees, as long as the test data is exchangeable with the training data. Complex scenarios involving multiple predictions can therefore result in multiple conformal prediction sets—for example, when there are multiple trials, models, or data sources.
@@
Aggregating multiple prediction sets provides a summary view of the uncertainty across predictions. Various aggregation methods exist, each with their own balance of sample efficiency and computational cost. For conformal prediction, where the popular split conformal variant [14] introduces a one-time random split, there are multiple methods of aggregating predictions to reduce the randomness over multiple splits: popular examples include cross-conformal [15], CV+ [16], and jackknife+ [16]. These aggregation methods are all symmetric, in that individual prediction sets contribute equally to the aggregate set. Although comparatively less studied, it is often useful to aggregate sets asymmetrically: that is, to weight sets based on their prior importance to the overall result [17]. Work in both symmetric and asymmetric aggregation establishes that aggregating individual prediction sets with $1-\alpha$ coverage guarantees results in an overall coverage guarantee of $1-2\alpha$ [18].
@@
Our work is based on the observation that if the overall coverage guarantee for asymmetric aggregation reflects the contributions of the individual sets, then we can achieve a tighter guarantee than a constant $1-2\alpha$. Consider, as an extreme, the case where almost all the weight is on one set; then, because the coverage of that set dominates the aggregation, the overall coverage guarantee should be closer to the $1-\alpha$ guarantee of the dominant individual set. To this end, we propose a method for asymmetric aggregation where the overall coverage guarantee is based on the distribution of weights across the prediction sets. This approach leverages the results of Vovk and Wang [18] on averaging p-values to achieve tighter guarantees when there exists strong asymmetry in the importance of the prediction sets—i.e. when a single prediction set is significantly more important than the rest—with $1-2\alpha$ as the worst-case guarantee when the contributions of the prediction sets are more equal.
@@
The results on averaging p-values require that the weights be independent of the prediction sets, and therefore of the data. However, it is often desirable for weights to reflect information about the data—such as in the popular mixture-of-experts setting [19], where expert weights are learned from data. To deal with this general case, we additionally propose a method for aggregating prediction sets with data-dependent weights via a linear transformation on the weighted average of p-values associated with each set. Our method allows us to construct a valid aggregate prediction set while also preserving the proportions of the weights assigned to the individual sets. Since data-dependent weights adjust the contributions of the individual sets based on the observed input, our method also achieves a form of conditional/data-adaptive coverage. We apply our method to the mixture-of-experts setting, and we demonstrate its effectiveness in experiments with real and synthetic data.
@@
  • We propose a framework for the asymmetric aggregation of prediction sets based on weighted p-values (§4). With data-independent weights, our framework improves the coverage guarantee beyond the standard $1-2\alpha$, depending on the distribution of weights (Prop. 4.1).
  • We derive a general method for transforming a random variable to a p-value variable, which enables the construction of a prediction set with finite-sample guarantees (§5). Our method allows us to extend our framework to data-dependent weights, and we demonstrate that incorporating data-dependent weights provides a form of (locally) conditional coverage (Prop. 5.1, Prop. 5.2).
@@ 2 Background
Quantile and the empirical quantile function   Given a random variable $X\in\mathcal{X}$, the left $\alpha$-quantile of $X$, which we refer to as the $\alpha$-quantile of $X$ for simplicity, is defined as
@@
where $-Z=\{-z_{1},\dots,-z_{n}\}$. Note that the $n+1$ term serves as a finite sample correction.
p-values   A p-variable [18] is a random variable $P^{\prime}$ such that
@@
$\widehat{p}(x,y)=\frac{1+\sum_{i\in S_{\text{cal}}}\mathbbm{1}\{s(x,y)<s(X_{i},Y_{i})\}}{|S_{\text{cal}}|+1},$
which returns the proportion of calibration points that are less conformal than some test point $(x,y)$, with a finite sample correction. (Note that for an infinite number of samples, this quantity is uniform under exchangeability.) The equivalent prediction set constructed from the p-value function is
@@
Since we only need to fit a single predictor $\widehat{\mu}$, split conformal is computationally cheap. However, it has a statistical cost: if $|S_{\text{cal}}|$ is small, then the variability/randomness in the process used to generate the prediction set is high.
@@ 3 Related work
The prediction set for a test point depends on both the data generating process and the choice of model $\widehat{\mu}$; in split conformal, it also depends on how the data is split. This last dependence we can mitigate by repeatedly performing the split conformal procedure and combining the results. In this section, we categorize existing work on combining conformal results into two main approaches: the first focuses on methods for selecting or combining prediction sets, while the second focuses on methods for combining p-values derived from multiple runs of the conformal procedure.
@@
Rather than combine prediction sets, Yang and Kuchibhotla [21] propose to select the predictor, from a set of $K$ predictors, that returns the best prediction set. They present two methods: one to select the predictor that gives the most efficient prediction set, but with only approximate validity, and one to select the predictor with the most valid prediction set (with minimal coverage slack), albeit with a width only close to the minimum. Liang et al. [22] point out that the approximate validity of the first method is due to selection bias, and introduce an alternative approach that both selects the most efficient predictor and uses it to construct a valid prediction set.
A recent line of work explores combining conformal prediction sets by majority vote with a $1-2\alpha$ coverage guarantee [23, 24, 17]. In particular, Gasparin and Ramdas [17] characterize the width of the majority vote set and introduce many extensions, including a weighted majority vote method that incorporates prior information in the weights. Our work on combining conformal results by weighted p-values is closely related to this method. However, our formulation allows us to leverage the results of Vovk and Wang [18] to improve the coverage guarantee beyond the standard $1-2\alpha$.
@@
A substantial body of research has been devoted to developing methods for combining p-values. Here, we focus specifically on approaches that accommodate arbitrary dependence, with a particular emphasis on their applications to conformal prediction. (For a more comprehensive review, see Balasubramanian et al. [25] and DiCiccio et al. [26].) An early example of combining p-values with arbitrary dependence is the Bonferroni method, where the minimum of a set of p-values is scaled by the number of p-values. This method has multiple extensions [27, 28], and was first applied to conformal prediction by Lei et al. [20]. Rüschendorf [29] finds that twice the average of p-values is a p-value; this was later extended to a more general notion of average by Vovk and Wang [18]. Stutz et al. [30] use the result of Rüschendorf [29] to get a $1-2\alpha$ guarantee for the average of p-values, but also propose a novel transformation to get a p-value average with improved coverage. Liu and Xie [31] propose the Cauchy combination test, where the test statistic is a weighted sum of Cauchy-transformed p-values. Wu et al. [32] apply the Cauchy combination test to conformal prediction, and show that their method yields asymptotically exact prediction sets under a weak stability assumption and a joint distribution assumption.
@@ 4 Combining conformal prediction sets by weighted p-value
$\widehat{p}_{k}(x,y)=\frac{1+\sum_{i\in S_{k}}\mathbbm{1}\{s_{k}(x,y)<s_{k}(X_{i},Y_{i})\}}{|S_{k}|+1}.$ (2)
@@
Now, consider assigning a weight $v_{k}$ to each p-value function $\widehat{p}_{k}$. The weighted average p-value function is then
@@
We now provide a coverage guarantee for this aggregated prediction set.
Proposition 4.1.
Let $\widehat{C}_{\alpha,1}^{\text{split}}(X_{n+1}),\dots,\widehat{C}_{\alpha,K}^{\text{split}}(X_{n+1})$ be $K$ prediction sets defined by p-value functions $\widehat{p}_{1},\dots,\widehat{p}_{K}$ (2) on $X_{n+1}$, where $1-\alpha$ coverage (3) holds for each set $k\in[K]$. Then, the prediction set $\widehat{C}_{\alpha}^{\text{avg}}(X_{n+1})$ (5) defined by the weighted average p-value function (4) has the following coverage guarantee:
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{avg}}_{\alpha}(X_{n+1})\right\}\geq 1-\min\left\{\frac{1}{v},2\right\}\alpha\quad\forall\alpha\in(0,1),$ (6)
@@
In a sense, this result provides a spectrum of coverage guarantees based on the weight distribution among the models. If one predictor dominates (i.e., $v>1/2$), the coverage guarantee improves beyond the standard $1-2\alpha$ of typical aggregation methods, allowing the overall coverage to approach that of the most influential predictor. In the extreme case, when all the weight is assigned to a single model ($v=1$), we recover the standard split conformal guarantee of $1-\alpha$ coverage for that model. Thus, we can interpolate between the standard $1-2\alpha$ guarantee of combined models and the $1-\alpha$ guarantee of individual models, with the weights controlling the trade-off. This improves the coverage guarantee for asymmetric aggregation, and opens up the utility of asymmetric generalizations of the many established symmetric aggregation methods (e.g. set-weighted versions of cross-conformal, CV+, etc.)
Since the work of Vovk and Wang holds for arbitrarily dependent p-values, this method is robust across a wide range of scenarios. In practice, this weighting scheme allows users to incorporate prior knowledge about the relative quality of different predictors [18, 17]. The weighting can reflect, for example, expert insights on which of $K$ models should be prioritized as being more reliable.
@@ 5 Extending to data-dependent weights
In many scenarios, it is desirable to assign model weights based on the observed data. Data-dependent weights offer an adaptive approach that adjusts the influence of each model according to specific characteristics of the data. However, when weights depend on the data, they also become dependent on the p-variables; as a result, the theory of Vovk and Wang [18] does not directly apply to this setting. Nevertheless, we can build on their approach to propose a simple, empirical method that allows us to recover a coverage guarantee, even when the weights depend on the data.
@@
Let the p-variables of the predictors be $P_{1},\dots,P_{K}\in\mathcal{U}$, where $\mathcal{U}$ is the set of all uniformly distributed random variables. The weights are given by a random vector $\mathbf{W}=(W_{1},\dots,W_{K})$ in the $(K-1)$-dimensional simplex $\Delta_{K-1}:=\{\mathbf{w}=(w_{1},\dots,w_{K})\in[0,1]^{K}:w_{1}+\dots+w_{K}=1\}$ depending on the data $\{(X_{i},Y_{i})\}_{i\in[n]}\cup\{X_{n+1}\}$. We define the weighted average function
@@
(See §A.4 for a full derivation.) The scaling defined in (8) transforms the weighted average $P_{\text{all}}$ into a valid p-variable, recovering a coverage guarantee. We formalize this in the following result.
Proposition 5.1.
Let $\widehat{C}_{\alpha,1}^{\text{split}}(X_{n+1}),\dots,\widehat{C}_{\alpha,K}^{\text{split}}(X_{n+1})$ be $K$ prediction sets defined by p-value functions $\widehat{p}_{1},\dots,\widehat{p}_{K}$ (2) on $X_{n+1}$ corresponding to p-variables $P_{1},\dots,P_{K}$. Suppose $1-\alpha$ coverage (3) holds for each set $k\in[K]$. Let $(W_{1},\dots,W_{K})$ be a random vector in $\Delta_{K-1}$ depending on $\{(X_{i},Y_{i})\}_{i\in[n]}\cup\{X_{n+1}\}$. Then, the prediction set $\widehat{C}_{\alpha}^{\text{data-avg}}(X_{n+1})$ defined by the weighted average $p_{\text{all}}$ (7) corresponding to random variable $P_{\text{all}}$ has the following coverage guarantee:
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\right\}\geq 1-m^{*}\alpha\quad\forall\alpha\in(0,1),$ (9)
where $m^{*}$ is defined in (8).
@@
1:Input: $K$ predictors from potentially overlapping datasets, $K$ weights dependent on the predictors (i.e. potentially learned from the same data), data with indices $\{1,\dots,n\}=S_{\text{cal}}\cup S_{\text{merge}}$, and test example $X_{n+1}$.
2:Output: Prediction set around the test example $\widehat{C}_{\alpha}^{\text{data-avg}}(X_{n+1})$ with coverage $1-\widehat{m}^{*}\alpha$.
@@
5: 1.1. Using calibration set $S_{\text{cal}}$, derive p-value functions $\widehat{p}_{k}$ for each of the predictors (2).
6: 1.2. Take the weighted sum of the p-value functions $\widehat{p}_{k}$ with the $K$ dependent weights. The result is the function $p_{\text{all}}$, with prediction set $\widehat{C}_{\alpha}^{\text{data-avg}}(X_{n+1})$.
@@
10: 2.2. Derive the correction factor $\widehat{m}^{*}$ from the samples of $\widehat{F}_{P_{\text{all}}}$ (11).
@@
The exact value of $m^{*}$ depends on $F_{P_{\text{all}}}$, which is generally unknown. We therefore propose to set aside a subset of the data $S_{\text{merge}}$ [^2] So named because $S_{\text{merge}}$ is used to learn the correction that makes the merging function [18] for the weighted average empirically precise. to estimate $F_{P_{\text{all}}}$ in practice.
Typically, an $n$-sample empirical CDF for $P_{\text{all}}$ would be given by
$\widehat{F}_{P_{\text{all}}}(\alpha)=\frac{\sum_{i\in\mathcal{S}_{\text{merge}}}\mathbbm{1}\{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})\leq\alpha\}}{|\mathcal{S}_{\text{merge}}|};$
we use the more conservative function
@@
as it tends to yield more stable estimates in finite samples. The correction factor is then computed as
$\widehat{m}^{*}=\max_{i\in\mathcal{S}_{\text{merge}}}\frac{\widehat{F}^{\text{cons}}_{P_{\text{all}}}(p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)}))}{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})},$ (11)
following (8). (For more detail on why (11) is equivalent to (8) for an empirical CDF, see §A.5.) The computation of $\widehat{m}^{*}$ yields the following finite-sample coverage guarantee:
Proposition 5.2.
(Finite-sample guarantee.) Under the same assumptions as Proposition 5.1, fix a user-chosen failure probability $\delta\in(0,1)$ and set

### v2

@@ Abstract
Conformal prediction quantifies the uncertainty of machine learning models by augmenting point predictions with valid prediction sets. For complex scenarios involving multiple trials, models, or data sources, conformal prediction sets can be aggregated to create a prediction set that captures the overall uncertainty, often improving precision. However, aggregating multiple prediction sets with individual $1-\alpha$ coverage inevitably weakens the overall guarantee, typically resulting in $1-2\alpha$ worst-case coverage. In this work, we propose a framework for the weighted aggregation of prediction sets, where weights are assigned to each prediction set based on their contribution. Our framework offers flexible control over how the sets are aggregated, achieving tighter coverage bounds that interpolate between the $1-2\alpha$ guarantee of the combined models and the $1-\alpha$ guarantee of an individual model depending on the distribution of weights. Importantly, our framework generalizes to data-dependent weights, as we derive a procedure for weighted aggregation that maintains finite-sample validity even when the weights depend on the data. This extension makes our framework broadly applicable to settings where weights are learned, such as mixture-of-experts (MoE), and we demonstrate through experiments in the MoE setting that our methods achieve adaptive coverage.
@@ 1 Introduction
In recent years, machine learning models have achieved remarkable accuracy across a wide range of predictive tasks [1, 2, 3]. Understanding the uncertainty associated with each prediction is essential for decision-making in real-world scenarios, but the black-box nature of many machine learning models hinders their deployment in safety-critical applications such as medical diagnosis [4, 5, 6, 7], industrial control systems [8, 9], and extreme weather forecasting [10, 11, 12]. Conformal prediction [13] emerged as a popular wrapper method around machine learning models because it provides a statistically valid quantification of uncertainty, transforming point predictions to prediction sets with finite-sample coverage guarantees, as long as the test data is exchangeable with the training data.
@@
Complex scenarios with multiple predictions—for example, when there are multiple trials, models, or data sources—naturally produce multiple conformal prediction sets (Figure 1, left). A number of methods have been proposed to aggregate prediction sets. For conformal prediction, where the popular split conformal variant [14] introduces a one-time random split, there are multiple methods of aggregating predictions to reduce the randomness over multiple splits: popular examples include cross-conformal [15], CV+ [16], and jackknife+ [16]. These aggregation methods are all symmetric, in that individual prediction sets contribute equally to the aggregate set. Although comparatively less studied, a more general approach is to aggregate sets asymmetrically—that is, to weight sets based on their prior importance to the overall result [17]. Work in both symmetric and asymmetric aggregation establishes that aggregating individual prediction sets with $1-\alpha$ coverage guarantees results in an overall coverage guarantee of $1-2\alpha$ [18].
Example: storm forecasting Refer to caption Refer to caption
(a) Weights are data-independent4)
  • tighter guarantee than previous work (Prop. 5)
  • expert priors: “humidity is generally more reliable”
(b) Weights are data-dependent5)
  • adaptive coverage (e.g. mixture-of-experts §6)
  • personalized predictions: “for $X$ location, wind shear is more important than sea temperature”
Figure 1: Left: Storm forecasting example with different models tracking humidity, sea temperature, rainfall, and wind shear over time. Below, an abstract representation of how models vary in predictive strength across the input space (colored regions). At the given test point (black), the red and green models dominate, so their prediction sets matter most. Right: Model prediction sets are combined with weighted aggregation. (a) Data-independent weights reflect expert priors (e.g. the blue model covers the largest portion of the input space and so is considered more generally reliable, and up-weighted accordingly). (b) Data-dependent weights adapt to context, yielding forecasts better aligned with current conditions (e.g. at the test point, red and green dominate).
@@
Our work is based on the observation that if the overall coverage guarantee for asymmetric aggregation reflects the contributions of the individual sets, then we can achieve a tighter guarantee than a constant $1-2\alpha$. Consider, as an extreme, the case where almost all the weight is on one set; then, because the coverage of that set dominates aggregation, the overall coverage guarantee should be closer to the $1-\alpha$ guarantee of the dominant individual set. Accordingly, we propose a method for asymmetric aggregation with data-independent weights where the overall coverage guarantee is based on the distribution of weights across the prediction sets. This approach leverages the results of Vovk and Wang [18] on averaging p-values to achieve tighter guarantees when there exists strong asymmetry in the importance of the prediction sets—i.e. when a single prediction set is significantly more important than the rest—with $1-2\alpha$ as the worst-case guarantee when the contributions of the prediction sets are more equal.
@@
Data-independent weights can encode expert priors on the relative importance of the prediction sets, allowing weights to reflect prior information such as the reliability of the predictors used to generate the sets (Figure 1a). Already, these prior weights generalize standard symmetric aggregation. Yet in practice, many applications demand even more flexibility, and require weights that adapt directly to the data: for example, in the popular mixture-of-experts setting [19], expert weights are learned from inputs. To deal with this general case, we additionally propose a method for aggregating prediction sets with data-dependent weights via a linear transformation on the weighted average of p-values associated with each set. Our method allows us to construct a valid aggregate prediction set while also preserving the proportions of the weights assigned to the individual sets. Since data-dependent weights adjust the contributions of the individual sets based on the observed input, our method also achieves adaptive coverage (Figure 1b). We apply our method to the mixture-of-experts setting, and we demonstrate its effectiveness in experiments with real and synthetic data.
@@
  • We propose a framework for the asymmetric aggregation of prediction sets based on weighted p-values (§4). With data-independent weights, our framework improves the coverage guarantee beyond the standard $1-2\alpha$, depending on the distribution of weights (Prop. 5).
  • We derive a general method for transforming a random variable to a p-value variable, which enables the construction of a prediction set with finite-sample guarantees (§5). This method allows us to extend our framework to data-dependent weights, and we demonstrate that data-dependent weights provide adaptive coverage (Prop. 5, Prop. 5).
@@ 2 Background
Quantile and the empirical quantile function
Given a random variable $X\in\mathcal{X}$, the left $\alpha$-quantile of $X$, which we refer to as the $\alpha$-quantile of $X$ for simplicity, is defined as
@@
where $Z=\{-z_{1},\dots,-z_{n}\}$. Note that the $n+1$ term serves as a finite sample correction.
p-values
A p-variable [18] is a random variable $P^{\prime}$ such that
@@
$\widehat{p}(x,y)=\frac{1+\sum_{i\in S_{\text{cal}}}\mathds{1}\{s(x,y)<s(X_{i},Y_{i})\}}{|S_{\text{cal}}|+1},$
which returns the proportion of calibration points that are less conformal than some test point $(x,y)$, with a finite sample correction. The equivalent prediction set constructed from the p-value function is
@@ 3 Related work
We categorize existing work on combining conformal results into two approaches: methods for selecting or combining prediction sets, and methods for combining p-values derived from multiple runs of the conformal procedure.
@@
Rather than combine prediction sets, Yang and Kuchibhotla [21] propose to select the predictor, from a set of $K$ predictors, that returns the best prediction set. They present two methods: one to select the predictor that gives the most efficient prediction set, but with only approximate validity, and one to select the predictor with the most valid prediction set (i.e. with minimal coverage slack), albeit with a width only close to the minimum. Liang et al. [22] point out that the approximate validity of the first method is due to selection bias, and introduce an alternative approach that both selects the most efficient predictor and uses it to construct a valid prediction set.
A recent line of work explores combining conformal prediction sets by majority vote with a $1-2\alpha$ coverage guarantee [23, 24, 17]. In particular, Gasparin and Ramdas [17] characterize the width of the majority vote set and introduce many extensions, including a weighted majority vote method that incorporates prior information in the weights. Our work on combining conformal results by weighted p-values is closely related to this method. However, our formulation allows us to leverage the results of Vovk and Wang [18] to improve the coverage guarantee beyond the standard $1-2\alpha$, and can also be generalized to data-dependent weights to enable adaptive coverage. (See §C.1 for more detail.)
@@
A substantial body of research has been devoted to developing methods for combining p-values. Here, we focus specifically on approaches that accommodate arbitrary dependence, with a particular emphasis on their applications to conformal prediction. (For a more comprehensive review, see Balasubramanian et al. [25] and DiCiccio et al. [26].)
An early example of combining p-values with arbitrary dependence is the Bonferroni method, where the minimum of a set of p-values is scaled by the number of p-values. This method has multiple extensions [27, 28], and was first applied to conformal prediction by Lei et al. [20].
Rüschendorf [29] finds that twice the average of p-values is a p-value, a result that was later extended to a more general notion of average by Vovk and Wang [18]. Stutz et al. [30] use the result of Rüschendorf [29] to get a $1-2\alpha$ guarantee for the average of p-values, but also propose a novel transformation to get a p-value average with improved coverage. (For a discussion of how to extend their method to apply to our framework, see §C.2.)
Liu and Xie [31] propose the Cauchy combination test, where the test statistic is a weighted sum of Cauchy-transformed p-values. Wu et al. [32] apply the Cauchy combination test to conformal prediction, and show that their method yields asymptotically exact prediction sets under a weak stability assumption and a joint distribution assumption.
@@ 4 Combining conformal prediction sets by weighted p-value
$\widehat{p}_{k}(x,y)=\frac{1+\sum_{i\in S_{k}}\mathds{1}\{s_{k}(x,y)<s_{k}(X_{i},Y_{i})\}}{|S_{k}|+1}.$ (2)
@@
Consider assigning a weight $v_{k}$ to each p-value function $\widehat{p}_{k}$. The weighted average p-value function is then
@@
We now provide a coverage guarantee for this aggregated prediction set. {restatable}propositionCoverageConstantW Let $\widehat{C}_{\alpha,1}^{\text{split}}(X_{n+1}),\dots,\widehat{C}_{\alpha,K}^{\text{split}}(X_{n+1})$ be $K$ prediction sets defined by p-value functions $\widehat{p}_{1},\dots,\widehat{p}_{K}$ (2) on $X_{n+1}$, where $1-\alpha$ coverage (3) holds for each set $k\in[K]$. Then, the prediction set $\widehat{C}_{\alpha}^{\text{avg}}(X_{n+1})$ (5) from thresholding the weighted average p-value function (4) gives the coverage guarantee
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{avg}}_{\alpha}(X_{n+1})\right\}\geq 1-\min\left\{\frac{1}{v},2\right\}\alpha\quad\forall\alpha\in(0,1),$
@@
In a sense, this result provides a spectrum of coverage guarantees based on the weight distribution among the models. When one predictor dominates (i.e., $v>1/2$), the guarantee improves beyond the standard $1-2\alpha$ of typical aggregation methods, allowing the overall guarantee to approach that of the most influential predictor. In the extreme case, when all the weight is assigned to a single model ($v=1$), we recover the standard split conformal guarantee of $1-\alpha$ coverage for that model. Thus, we can interpolate between the standard $1-2\alpha$ guarantee of combined models and the $1-\alpha$ guarantee of individual models, with the weights controlling the trade-off. This improves the coverage guarantee for asymmetric aggregation, and also opens up the utility of asymmetric generalizations of the many established symmetric aggregation methods (e.g. set-weighted versions of cross-conformal, CV+, etc.)
Since the work of Vovk and Wang holds for arbitrarily dependent p-values, this method is robust across a wide range of scenarios. In practice, independent weights allow users to incorporate prior knowledge about the relative quality of different predictors [18, 17]. The weighting can reflect, for example, expert insights on which of $K$ models should be prioritized as being more reliable.
@@ 5 Extending to data-dependent weights
In many practical settings, model weights are determined by the observed data rather than fixed in advance. Such data-dependent weights naturally adapt the influence of each model to the characteristics of the input, making them central to applications like ensemble learning and mixture-of-experts. This adaptivity, however, creates a technical challenge: once the weights depend on the data, they also depend on the associated p-variables, and the theory of Vovk and Wang [18] no longer applies. Nevertheless, we can build on their approach to develop a method that allows us to recover a valid coverage guarantee, even when the weights are data-dependent.
@@
Let the p-variables of the predictors be $P_{1},\dots,P_{K}\in\mathcal{U}$, where $\mathcal{U}$ is the set of all uniformly distributed random variables. The weights are given by a random vector $\mathbf{W}=(W_{1},\dots,W_{K})$ in the (K1)(K-1)-dimensional simplex $\Delta_{K-1}:=\{\mathbf{w}=(w_{1},\dots,w_{K})\in[0,1]^{K}:w_{1}+\dots+w_{K}=1\}$ depending on the data $\{(X_{i},Y_{i})\}_{i\in[n]}\cup\{X_{n+1}\}$. We define the weighted average function
@@
(See §A.4 for a derivation of the supremum expression.) The scaling defined in (7) transforms the weighted average $P_{\text{all}}$ into a valid p-variable, recovering a coverage guarantee. We formalize this in the following result.
{restatable}
[Infinite-sample guarantee]propositionCoverageDataW Let $\widehat{C}_{\alpha,1}^{\text{split}}(X_{n+1}),\dots,\widehat{C}_{\alpha,K}^{\text{split}}(X_{n+1})$ be $K$ prediction sets defined by p-value functions $\widehat{p}_{1},\dots,\widehat{p}_{K}$ (2) on $X_{n+1}$ corresponding to p-variables $P_{1},\dots,P_{K}$. Suppose $1-\alpha$ coverage (3) holds for each set $k\in[K]$. Let $(W_{1},\dots,W_{K})$ be a random vector in $\Delta_{K-1}$ depending on $\{(X_{i},Y_{i})\}_{i\in[n]}\cup\{X_{n+1}\}$, and let $p_{\text{all}}$ be the weighted average (6), with random variable $P_{\text{all}}$.
The prediction set from thresholding $p_{\text{all}}$ is $\widehat{C}_{\alpha}^{\text{unscaled}}(X_{n+1})=\{y:p_{\text{all}}(X_{n+1},y)>\alpha\}$, and it satisfies the coverage guarantee
$\mathbb{P}\{Y_{n+1}\in\widehat{C}^{\text{unscaled}}_{\alpha}(X_{n+1})\}\geq 1-m^{*}\alpha\quad\forall\alpha\in(0,1).$
The prediction set from thresholding $m^{*}p_{\text{all}}$ is $\widehat{C}_{\alpha}^{\text{scaled}}(X_{n+1})=\{y:m^{*}p_{\text{all}}(X_{n+1},y)>\alpha\}$, and it satisfies the coverage guarantee
$\mathbb{P}\{Y_{n+1}\in\widehat{C}^{\text{scaled}}_{\alpha}(X_{n+1})\}\geq 1-\alpha\quad\forall\alpha\in(0,1).$
Here, $m^{*}$ is the correction factor defined in (7).
@@
1:Input: Data with indices $\{1,\dots,n\}=S_{\text{cal}}\cup S_{\text{merge}}$, $K$ predictors from potentially overlapping datasets, $K$ weights which may depend on data (e.g. learned router weights), test example $X_{n+1}$.
2:Output: Prediction set around the test example $\widehat{C}_{\alpha}^{\text{scaled}}(X_{n+1})$, with coverage of at least $1-(\alpha+\epsilon+\delta)$.
@@
5: $1.$ Using calibration set $S_{\text{cal}}$, derive p-value functions $\widehat{p}_{k}$ for each of the predictors (2).
6: 1.2. Compute the aggregated p-value function $p_{\text{all}}=\sum_{k=1}^{K}w_{k}\,\widehat{p}_{k}$, using the (potentially data-dependent) weights $w_{k}$.
@@
10: $2.$ Derive the correction factor $\widehat{m}^{*}$ from the samples of $\widehat{F}_{P_{\text{all}}}$ (9).
11:
12:Step 3. For test example $X_{n+1}$, construct the prediction set
$\widehat{C}^{\text{scaled}}_{\alpha}(X_{n+1})=\{y:\widehat{m}^{*}p_{\text{all}}(X_{n+1},y)>\alpha\}.$
@@
By definition, $m^{*}$ is the minimal scaling factor that makes $P_{\text{all}}$ a valid p-variable, and its value is determined by the CDF $F_{P_{\text{all}}}$. Because $F_{P_{\text{all}}}$ is a population quantity that is not accessible in practice, we cannot evaluate $m^{*}$ directly. Instead, we define a computable proxy $\widehat{m}^{*}$ using an empirical CDF constructed from a designated merging set $S_{\text{merge}}$ [^2] So named because $S_{\text{merge}}$ is used to learn the correction that makes the merging function [18] for the weighted average empirically precise..
Typically, the empirical CDF for $P_{\text{all}}$ computed from $S_{\text{merge}}$ is given by
$\widehat{F}_{P_{\text{all}}}(\alpha)=\frac{\sum_{i\in S_{\text{merge}}}\mathds{1}\{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})\leq\alpha\}}{|S_{\text{merge}}|},$
but in practice we use the more conservative estimate
@@
as it tends to yield more stable estimates in finite samples. The corresponding empirical correction factor is
$\widehat{m}^{*}=\max_{i\in S_{\text{merge}}}\frac{\widehat{F}^{\text{cons}}_{P_{\text{all}}}(p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)}))}{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})},$ (9)
following (7). (For more detail on why (9) is equivalent to (7) for an empirical CDF, see §A.5.) The construction of $\widehat{m}^{*}$ yields the following finite-sample guarantee:
{restatable}
[Finite-sample guarantee]propositionCoverageDataWEmpirical Under the same assumptions as Proposition 5, fix a user-chosen failure probability $\delta\in(0,1)$ and set